# ReCAP: Recursive Cross Attention Network for Pseudo-Label Generation in Robotic Surgical Skill Assessment

### v2

@@ IV-A Performance
To the best of our knowledge, we are the only work to report performance metrics on the JigSaw Dataset in a task-agnostic manner. Recent studies have focused on video-centric models, which don’t generalise well outside of their training domain, whereas pure kinematic, a task-agnostic modality, has seen decreased interest. This research showcases that there is still substantial room for improvement in the kinematic domain on the JigSaw dataset. This work showcased competitive performance on the GRS. It doesn’t beat SOA in the video domain, certainly due to overfitting. The ratio between input features and model parameters will always lead to overfitting in deep-learning applications (238644 / 1440). On the OSATS, our model performs better than most, and we thoroughly report on our performance. We do see that for Overall Performance in Needle Passing and Quality of the Final Product across the three tasks, the model doesn’t perform as well. We argue that this is because while kinematic data is an extremely powerful modality it cannot capture all the nuances in performance. It was shown that the quality-of-final-product by Kasa et al. [21] is an image/video-centric score. Furthermore, it is also worth noting that similar kinematic profiles might end up with very different performance scores. In needle passing if a clinician misgauges the depth, he will fail the task, and receive a low Overall Performance score which wouldn’t necessarily be translated in their kinematic profile. As highlighted by Lefor et al. [29], the JIGSAW dataset is very imbalanced and inverse correlation among subjects can be observed. Furthermore using Speaman’s correlation coefficient when there are sometimes only 3 test samples could be misleading. Consequently, developing tools based on this dataset may not seamlessly translate to real-world practice. We argue that our competitive results, despite a simple architecture, are because our problem formulation allows the model to attend segments of the input independently, while retaining some temporal context, allowing for much more flexible behaviour and sparse solutions.
IV-B Proposed Method
@@
Through the ablations, we showcase that using the pseudo-label in the loss boosts the model’s performance. We hypothesise that including intermediate predictions prevents overfitting and acts as a regularization component. Furthermore this also fits within how raters would look at performance, where each previous segment affects their decision. However the way the loss is built wouldn’t allow for the model to understand when a catastrophic event would occur, which would always result in a systematic low performance score. Allowing the model to assign a weight to the segments should adress this issue.
IV-C Pseudo-Labels
@@
The contribution of this loss is two-fold, improving performance and generating pseudo-labels. We extended the generation of pseudo-labels to insights into the user’s performance. We can visualize our pipeline results in Fig. 2. As the introduction emphasised, mapping performance to actionable feedback is necessary for clinicians’ understanding and the first steps toward clinical translation of such models. It is also worth noting the recurrent nature of the model allows for its online application. However, the validation of those pseudo-labels has serious limitations. Distinguishing between good/ average and average/poor levels at such a fine-grained scale presents significant complexity for any rater. This is obvious when looking at the agreement rate of the surgeon with random noise. The expected agreement rate should’ve been around 33%. Additionally, variability among raters is commonly encountered in various fields and tasks, making the extraction of ground truth complex. In an ideal scenario, having more data and involving multiple raters could help mitigate this variability. In practice, this is not feasible, and methods to extract labels in a weakly-supervised might be a more promising approach. The significant difference between random noise and the model’s prediction showcases a promising first step towards that goal, for both performance gains and interpretability.
@@ V Conclusion and Future Work
In this work, we adopt a different take on the problem formulation for skill assessment which can be expanded to more complex recurrent architectures and other fields of skill assessment. Through our competitive results on the JIGSAW dataset, we demonstrate the feasibility of this approach, while providing more granular performance insights for skill assessment. In future work, our focus will be on enhancing the robustness of validation, incorporating other time series data that an operating room could generate [33], and extending our approach to datasets featuring longer and more complex tasks. Surgical tasks/interventions lasting multiple hours are very demanding to annotate. Semi-supervised methods that extract intermediate labels, at the gesture, step, and phase levels could allow for more automated and more granular assessment of surgeons’ performance.

### v3

@@ IV-1 Performance
To the best of our knowledge, we are the only work to report performance metrics on the JIGSAWS dataset in a task-agnostic manner. Recent studies have focused on video-centric models, which don’t generalise well outside of their training domain, whereas pure kinematic, a task-agnostic modality, has seen decreased interest. This research showcases that there is still substantial room for improvement in the kinematic domain on the dataset. This work demonstrated competitive performance on the GRS, surpassing kinematic-only approaches and matching the performance of video-only methods, which are included for comparison and broader context. The ratio between input features and model parameters (238644 and 1440 respectively) will always lead to overfitting in deep-learning applications . Related to the performance capabilities of our model to predict OSATS, we rigorously evaluated our model, which overall outperformed on most tasks. However, we note that for Overall Performance in Needle Passing and Quality of the Final Product across the three tasks, the model doesn’t perform as well. We argue that this is because while kinematic data is an extremely powerful modality it cannot capture all the nuances in performance. It was shown that the quality-of-final-product by Kasa et al. [21] is an image/video-centric score. Furthermore, it is also worth noting that similar kinematic profiles might end up with very different performance scores. In needle passing if a clinician misgauges the depth, he will fail the task, and receive a low Overall Performance score which wouldn’t necessarily be translated in their kinematic profile. As highlighted by Lefor et al. [29], the JIGSAWS dataset is very imbalanced and inverse correlation between level of expertise and OSATS can be observed. Furthermore using Speaman’s correlation coefficient when there are sometimes only 3 test samples could be misleading. Consequently, developing tools based on this dataset may not seamlessly translate to real-world practice. We argue that our competitive results, despite a simple architecture, are because our problem formulation allows the model to attend segments of the input independently, while retaining some temporal context, allowing for much more flexible behaviour and sparse solutions.
IV-2 Proposed Method
@@
Through the ablations, we showcase that using the pseudo-label in the objective function boosts the model’s performance. We hypothesise that including intermediate predictions prevents overfitting and acts as a regularization component. Furthermore this also fits within how raters would look at performance, where each previous segment affects their decision. However the way the objective function is defined doesn’t allow for the model to understand when a catastrophic event would occur, which would always result in a systematic low performance score. Allowing the model to assign a weight to the segments should adress this.
IV-3 Pseudo-Labels
@@
The contribution of this loss is two-fold, improving performance and generating pseudo-labels. We extended the generation of pseudo-labels to insights into the user’s performance. We can visualize our pipeline results in Fig. 2. Following our motivation, mapping performance to actionable feedback is necessary for clinicians’ understanding and the first steps toward clinical translation of such models. It is also worth noting the recurrent nature of the model allows for its online application. However, the validation of those pseudo-labels has serious limitations. Distinguishing between good/ average and average/poor levels at such a fine-grained scale presents significant complexity for any rater. This is obvious when looking at the agreement rate of the surgeon with random noise. The expected agreement rate should’ve been around 33%. Additionally, variability among raters is commonly encountered in various fields and tasks, making the extraction of ground truth complex. In an ideal scenario, having more data and involving multiple raters could help mitigate this variability. In practice, this is not feasible, and methods to extract labels in a weakly-supervised might be a more promising approach. The significant difference between random noise and the model’s prediction showcases a promising first step towards that goal, for both performance gains and interpretability.
@@ V Conclusion and Future Work
In this work, we adopt a different take on the problem formulation for skill assessment which can be expanded to more complex recurrent architectures and other fields of skill assessment. Through our competitive results on the JIGSAWS dataset, we demonstrate the feasibility of this approach, while providing more granular performance insights for skill assessment. In future work, our focus will be on enhancing the robustness of validation, incorporating other time series data that an operating room could generate [33], and extending our approach to datasets featuring longer and more complex tasks. Surgical tasks/interventions lasting multiple hours are very demanding to annotate. Weakly-supervised methods that extract intermediate labels, at the gesture, step, and phase levels could allow for more automated and more granular assessment of surgeons’ performance.